OpenStudy (anonymous):

Solve for x: 2 over 5 (x - 2) = 4x

OpenStudy (anonymous):

whaaaat???????

OpenStudy (anonymous):

is the answer 9 or 2/9?

OpenStudy (anonymous):

condition \(\large x\neq2 \) .Fisrt, I multiply both sides by \(\large5 (x-2)\), get \(\large 2=4x\times5(x-2)\)\(\large \Rightarrow 2=4x\times5x-4x\times10\\\large \Rightarrow 20x^2-40x=2\\\large\Rightarrow 10x^2-20x-1=0\ \ (divide\ by\ 2)\) Then, using formula quadratic equation to find x:\(\large \color{blue}{x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}}\) Here \(\Large a=10, b=-20\ and\ c=-1\) Thus \(\Large x=\frac{20\pm\sqrt{20^2-4\times10\times(-1)}}{2\times10}\)

OpenStudy (texaschic101):

wait a second....is your problem : (2/5)(x - 2) = 4x ?? or 2/(5(x - 2)) = 4x ??

OpenStudy (texaschic101):

hello ?

OpenStudy (texaschic101):

JungHyunRan did it one way....here is the other way. You really need to make the problem clearer...add parenthesis. (2/5)(x - 2) = 4x 2/5x - 4/5 = 4x -- multiply by 5 2x - 4 = 20x -4 = 20x - 2x -4 = 18x -4/18 = x - 2/9 = x

OpenStudy (anonymous):

thank you so much. sorry for not replying I had to walk my dog

OpenStudy (anonymous):

it actually makes sense to me thanks to you @texaschic101

OpenStudy (texaschic101):

anytime :)